Journalist Report – January 07th

Journalist Report

Robinson Raphael, Crew 219 Astronomer,

Sol 2,

The Two Tales of the EVA

We are slowly adapting to our new life on Mars! The morning was filled with quiet as we were all tired from the past few days of repairs and moving in but slowly, we gained our strength after some coffee and a pleasant breakfast. A big chunk of the day consisted of two EVA’s. Meanwhile, the rest of the day afterward consisted of catching up on work around the Hab and personal time.

The first EVA started at 8am and it consisted of our CO, HSO, LSO, and GHO. The final destination was North Ridge and if time permitted, they had plans to climb it. Given the weather conditions, the crew proceeded with their journey and all things seemed well, right? Nope. A couple of the crewmembers had to deal with a lot fog in the helmets because of the cold! In the end, the crew did reach North Ridge but did not climb because of the fogging and snow cover. Despite all of it, the first EVA crew was rewarded with hot chocolate when they came back!

The second EVA started at 1pm and it consisted of myself, the XO, the engineer, and our geologist. The final destination was Kissing Camel Ridge E and plans were also made to climb to collect samples at different stratifications, if time permitted. This time around, the fogging issue had been solved and the weather was much better. Prior to heading to Kissing Camels, we set up the optical mount setup that I brought with me to MDRS. It was placed between the Musk Observatory and Robotic Observatory and it is marked with a flag that the Geologist and I made. Speaking of our Geologist, he was able to take some sweet photos of the optics mount. You can also see the mount from the MDRS habitat! A beautiful sight to see when people wake up if I do say so myself! After the mount was setup, we proceeded to drive the rovers to Kissing Camels.

Near Kissing Camel, we were able to collect some cool samples, each with their own interesting features. After collecting the samples, we saw that Kissing Camel had lots of area untouched by snow, so we proceeded to climb up the side of the ridge to collect additional samples. The climb went a bit slower than anticipated since it was my first time doing an EVA of this magnitude, but overall, we did what we came to do. In the end, the second EVA crew was also rewarded with hot chocolate, coffee, and an excellent lunch of tuna salad and pasta prepared by our LSO!

Operations Report – January 07th

Crew 219 Operations Report 07-JAN-2020

Sol: 2

Name of person filing report: Alejandro Luis Perez

Non-nominal systems: N/A

Generator: Run

Hours run: 13.5

From what time last night: 17:00

To what time this morning: 06:30

List of any additional daytime hours when the generator was run: N/A

Solar- SOC% (Before the generator is run at night): 49%

Diesel Reading – 74%

Station Propane Reading- 52%

Ethanol Free Gasoline- N/A

Water (loft tank) (gal) – 50

Water Meter (units)- 0147447,5

Water (static tank) (gal) – 526

Static to loft Pump used – Yes

Water in Green Hab (gal): 204.88

Water in Science Dome (gal): 0

Toilet Tank Emptied – No

Diemos rover used: Still in the Work shop

Hours:

Beginning Charge:

End Charge:

Currently Charging:

Sojourner rover used: Assigned to Director

Hours:

Beginning Charge:

End Charge:

Currently Charging:

Spirit Rover used- Yes

EVA #4

Starting Hours: 123.3

Beginning Charge: 100%

Ending Hours: 123.5

Ending Charge: 67%

Opportunity Rover used: Still in workshop

Hours:

Beginning Charge:

Ending Charge:

Currently Charging:

Curiosity Rover used: Yes

EVA #4

Starting Hours: 129.1

Beginning Charge: 100%

Ending Hours: 129.3

Ending Charge: 71 %

Notes on Rovers: Opportunity and Deimos off-site for maintenance.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): No

Reason for use:

Oil Added? No

# hours the ATV’s were used today:

Notes on ATVs:

Habcar used and why, where? Yes, used to fill static tank by Outpost.

CrewCar used and why, where? No

General Notes and Comments:

Summary of the internet: Nominal

Summary of Suit and Radios: Suits #1 and #7 are fully functional, weak air flow from yesterday’s EVA could have been from the low temperatures. Suit #5 does not work. The multimeter was used to see where power is not being transmitted and found that all connection from battery to circuit board are good. However, there is no power being delivered to the fans. Suit #3 has weak air flow coming from both fans even on full power. Battery testing on Suit #3 will be conducted tomorrow. Red Helmet was fixed by securing the head pad back in place using cable ties. The connector has been placed on the helmet missing a connector, however, there is no gasket for this connector. Also, one of the chargers got disconnected from its connector. This was fixed by soldering the cable back to the connector and this charger is working again.

Summary of Hab: Heating strip is working to keep the pipe from the static tank to loft tank from freezing. However, the insulation looks to be falling apart from the cold temperatures. Also, the cable obstructing the fire escape window has been removed and placed inside the RAM.

Summary of Science Dome operations: Nominal

Summary of Ram operations: Nominal

Summary of any observatory issues: Nominal

Summary of Health and safety issues: Nominal

Questions, concerns, and requests to Mission Support: Requesting Teflon tape for the loft tank feedline, new insulation for the Static tank to Loft tank pipe, and gasket for one of the 2 piece helmet connectors.

Research Report – January 07th

Crew 219 Science Report 07-JAN-20
Crew Science Officer: Hannah Blackburn

1. Biometrics and Neurobehavioral Research

Continued taking sleep and post-EVA logs. Post-EVA biometrics were recorded.

2. UAV

Nothing to report.

3. Dust Mitigation for Optical Mirrors

Optics mount and mirror were placed and marked with high-visibility flag.

4. Astrophotography of Celestial Bodies

Nothing to report.

5. Remediation of Mars Regolith

Regolith samples were placed in petri dishes and left to dry. Algae cells have been counted.

6. Chemical and Mineralogical Composition of the MDRS Site

Samples were taken from North Ridge and Kissing Camel on EVA 3 and 4. Samples were dried in the oven using glass beakers.

7. Protocols for the Discovery of Life on Mars

Samples were taken from North Ridge and Kissing Camel on EVA 3 and 4.

Glassware check out:

10 glass beakers (sizes between 200 and 500 ml) were used to dry regolith samples in the oven, and we will need them for the same purpose tomorrow. One glass pipette was used to measure BG11 media for the regolith remediation project.

Notes and Comments: There are unknown samples in the Science Dome incubator (which is not on) and have been there since the beginning of our sim. They are labeled “CowDung Road Gypsum 12/11” and appear to be a gel-like substance. Picture attached. Please advise on what to do with these samples.

GreenHab-Report – January 07th

Crew 219 GreenHab Report 07-JAN-20

Crew GreenHab Officer: Cynthia Montanez

Environmental control: Heating.

Shade cloth (40% and 30%) on.

Average temperature: 19.2 °C; 22.6 %

Average Max: 24.2 °C, 24.6%

Average Min: 14.3 °C, 18.6%

10:35 A.M.

Floor Unit: 11 °C

Electronic: 15.8 °C

humidity 23%

Max: 25.6 °C; 25%

Min: 10.9 °C; 17%

1:00 P.M.

Floor Unit: 13 °C

Electronic: 18.8 °C

humidity 21%

Max: 25.6 °C; 25%

Min: 19.5 °C; 20%

6:12 P.M.

Floor Unit: 9 °C

Electronic: 14.8 °C

humidity 23%

Max: 21.4 °C; 24%

Min: 12.7 °C; 19%

Hours of supplemental light: Light system 7:00PM-12:00AM.

Daily water usage for crops: ~8.37 Gal.

Daily water usage for research and/or other purposes: N/A

Water in Blue Tank – ~205.13 Gal.

Time(s) of watering for crops:

10:35 A.M.

SEASONAL PLANTING HISTORY:

Change to crops: None.

OTHER NOTES:

1st sprouts: N/A

Harvest: At 6:20 P.M. 4 grams of oregano and 3 grams of marjoram were harvested.

Other: The temperature in the Greenhab was colder than usual.

EVA Report 3 and 4 – January 07th

Crew 219 EVA Report 07-01-2020

EVA #3

Author: Dave Masaitis

Purpose of EVA: Geologic sample collection

Start time: 0903 hrs

End time: 1002 hrs

Narrative: EVA #3 entered the airlock at 0858 hours, and departed the Hab on foot at 0903. The EVA team moved north around the science dome and headed out towards North Ridge. Due to ambient temperature hovering around freezing, only a third of the distance towards North Ridge had been covered before the majority of crew members were reporting excessive visor fogging. CO called a short halt fogging and then resumed movement towards North Ridge. A second short halt had to be conducted just shy of the target sampling area for the same reason. Once on the target area at the base of North Ridge, crew members identified 5 locations to sample from in accordance with the crew geologist’s sample plan. The coring tool provided did a sufficient job of collecting the first sample, but the substrate was so frozen that the EVA team was unable to get the sample back out of the coring tool. CO crested a nearby ridge to reestablish communications with the Hab, and a secondary sampling technique was approved by the crew geologist. The EVA team collected and recorded locations of 10 samples, and then the CO and HSO discussed the plausibility of summiting North Ridge. Both concurred that between ground conditions and the persistent visor fogging being experienced by most of the team would make a summit attempt unsafe, so the EVA team opted to return to the Hab. EVA #3 reached the airlock at 1002 hours, for a total EVA duration of 59 minutes.

Destination: Vicinity of North Ridge

Coordinates: 0518600/4250800

Participants: D. Masaitis (CO), K. Crisman (HSO), H. Blackburn (LSO), C. Montanez (GHO)

Roads and routes per MDRS Map: North from the Hab towards North Ridge

Crew 219 EVA Report 07-01-2020

EVA #4

Author: Nathan Hadland

Purpose of EVA: Sample collection for mineralogy project and deployment of optical mount for dust mitigation project

Start time: 1250 hrs

End time: 1434 hrs

Narrative: EVA #4 initiated a 5-minute decompression at 1250 hours and moved out on foot to the vicinity of the Robotic Observatory. After setting up our experimental optical mount and safety flag to mark its location, we moved back to Curiosity and Spirit. After reporting SOC and hours, we progressed south along Cow Dung Road. We dismounted the rovers just past Robert’s Rock Garden and moved east on foot along the base of Kissing Camel Ridge E. The Crew Geologist identified a site with no snow cover and flat terrain for sample collection. We collected 5 samples in a grid and recorded GPS locations. After sample collection, we found a safe route up the side of Kissing Camel Ridge without snow cover and took 3 additional samples on the side of the formation. We then headed down and returned to the rovers. Curiosity and Spirit proceeded safely back to the Hab at 1428 hours and the EVA crew successfully pressurized in the air lock at 1434 hours.

Destination: Kissing Camel Ridge E

Coordinates: 0518100/4249300 (Kissing Camel Ridge E)

Participants: Nathan Hadland (XO), Abdul Elnagdi (GEO), Alejandro Perez (ENG), Robinson Raphael (ASTRO)

Roads and routes per MDRS Map: South along Cow Dung Road (0110), parked just south of Robert’s Rock Garden and then proceeded east on foot towards Kissing Camel Ridge E.

Mode of travel: Rovers (Spirit and Curiosity) and foot travel

Sol Summary – January 07

Crew 219 Sol Summary Report 07-01-2020

Sol: 2

Summary Title: Research, Research, Research

Author’s name: Dave Masaitis

Mission Status: Nominal

Sol Activity Summary: Sol 2 marks our full plunge into scientific research. With our mission plan published, we spent the day starting and continuing experiments, and conducted two more EVA’s for sample collection. Other features of the day include fresh bread and cobbler by the Lead Science Officer, and whispers of a showing of The Martian after the Comms window.

Look Ahead Plan: Tomorrow holds plans for two more proposed sample collection EVA’s, maintenance on a couple EVA suits, continued sample processing in the science dome, and continuation of our other experiments.

Anomalies in work: None

Weather: Cold and overcast all day, with temperatures below freezing.

Crew Physical Status: Nominal

EVA:

EVA #3 – Geologic Sample Collection near North Ridge (0518600/4250900)

EVA #4 – Geologic Sample Collection around Kissing Camels (0518100/4249300)

Reports to be filed:

Sol Summary
EVA Reports 1 & 2
EVA Requests 3 & 4
Operations Report
GreenHab Report
Research Report
Journalist’s Report
Photos (Including Photo of the Day)

Support Requested:

(Carried over from previous day)

Sundries: Paper towels
Medical: C-Spine needs to be replaces (broken/misshapen)
12x Medium Gauze Pads
Anitbiotic Ointment